# Geometric Model Selection for Latent Space Network Models: Hypothesis Testing via Multidimensional Scaling and Resampling Techniques

### v1

@@ 3.2 Network measures
The following network summary measures are used in this study:
@@ 3.3 Simulation details
To evaluate the efficiency of Method 1, we tested networks of size 15 to 75 with densities ranging from 0 to 0.9. To compare all three methods, we first tested networks of size 15 to 45 with densities ranging from 0 to 0.9. Then, we focus on large and sparse networks whose sizes ranges from 50 to 200 and densities up to 0.2.
@@ 3.4 Results
Efficiency of Method 1
As shown in Figure 2, we notice that as network sizes get larger, the hyperbolic networks tends to be correctly identified, and when network size is larger than 60, almost all hyperbolic networks are correctly categorized as hyperbolic. For smaller hyperbolic networks, the probability of choosing correct hyperbolic geometry gets larger as the density of the network gets larger. However, this method doesn’t work well for Euclidean networks, and larger and/or denser networks tends to be misclassified as hyperbolic network.
(a) Hyperbolic Networks
(b) Euclidean Networks
Figure 2: The Percentage of Networks that MDS Classified as Hyperbolic
Compare Three Methods
We adopt the language of diagnostic tests. Let sensitivity indicate how well a method correctly identifies hyperbolic networks, and specificity indicate how well a method correctly identifies Euclidean networks.
First, we focus on smaller networks with a wide range of densities.
N Density Method 1 Method 2 Method 3
@@
Table 1: Comparison of Three Proposed Methods on Smaller Networks
In general, for smaller networks, Method 1 could identify hyperbolic networks very well but, at the same time, tend to misclassify the majority of Euclidean networks as network getting larger and/or denser. Method 2 and Method 3 are more conservative and less likely to falsely categorize Euclidean networks as hyperbolic networks but misclassify the hyperbolic networks about more half the time.
Then, we focused on a particular group of networks - large and sparse networks. In this case, only networks of densities between 0 and 0.2 will be considered.
N Method 1 Method 2 Method 3
@@
Table 2: Comparison of Three Proposed Methods on Large and Sparse Networks
From Table 2, it is clear that as the networks become larger, when network size is larger than 100, Method 1 categorizes all simulated networks, hyperbolic and GLPM, as hyperbolic networks and loses the effect of distinguishing the underlying geometry. Both Method 2 and Method 3 show an improvement in distinguishing between two methods for large and sparse network, which is more common in the real-world datasets. Method 2 has better sensitivity, that is good at correctly identifying hyperbolic networks, while Method 3 has a better balance between sensitivity and specificity.
@@ 4.1 Karate Club
Papamichalis et al. [2022] use Zachary’s Karate club network as an example, compare the quality of MDS estimates when the manifold is assumed to be hyperbolic and Euclidean, and conclude that a model with a hyperbolic latent space better suits the data. Using the same package hydra as mentioned in this paper, we got $S_{\mathbb{R}^{2}}(\mathbf{Y})=24.65$ and $S_{\mathbb{H}^{2}}(\mathbf{Y})=18.20$. The stress from hyperbolic space is smaller, indicating a higher quality of embedding and shows that hyperbolic geometry would be better to model the data. Using Method 2 and Method 3, we get p-values of 0.0964 and 0.1468 respectively, which lead to the conclusion that Euclidean is better, as shown in Figure 3.
(a) Method 2, p-value: 0.0964
(b) Method 3, p-value: 0.1468
Figure 3: The Percentage of Networks that MDS Classified as Hyperbolic
@@ 4.2 Other Datasets
We have also applied these method on other datasets, a comparison of network characteristics and results from three different methods is presented below. The datasets used are from Nepusz et al. [2008](UKfaculty), Newman [2006](adjnoun), Lusseau et al. [2003](dolphin) and Kadushin [1995](ffe_friends).
(a) UKfaculty, n=81, dens.=0.178
Method 1: -19.32, Hyperbolic
Method 2: p-value=0, Hyperbolic
Method 3: N/A
(b) adjnoun, n=112, dens.=0.068
Method 1: -62.61, Hyperbolic
Method 2: p-value=0, Hyperbolic
Method 3: p-value=0, Hyperbolic
(c) dolphins, n=62, dens.=0.084
Method 1: 11.51, Euclidean
Method 2: p-value=0.99, Euclidean
Method 3: p-value=0.99, Euclidean
(d) ffe_friend, n=28, dens.=0.175
Method 1: -3.60, Hyperbolic
Method 2: p-value=0.20, Euclidean
Method 3: p-value=0.38, Euclidean
Figure 4: The average and standard deviation of critical parameters: Region R4
@@ 5 Discussion
We studied the effectiveness of using observed stress difference from MDS in determining the underlying geometry for latent space models through simulation study. It has been shown that when the stress difference is used alone, the geometry can’t be correctly identified especially for larger networks.
Two new methods based on the stress difference from MDS were proposed by introducing uncertainty with permutation test and bootstrapping. By studying the same sets of simulation networks, we showed that both methods outperform the original one, and work better on large and sparse networks which are commonly seen in real world.
The advantage of permutation test is that it retains the density of the observed network and doesn’t rely on the latent space models, but the structure of the network is not considered when generating the permuted networks. On the other hand, bootstrap method takes the structure of observed networks into account: if two nodes are currently connected, they are closer in latent space, thus more likely to be connected in the bootstrap networks. In this study, we used the Gaussian LPM as the Euclidean latent space model. However, for some networks, $\phi$ and $\tau$ could not be found through ad-hoc methods, thus bootstrapping could not be implemented.(Another thing about bootstrap method is that sometimes, it’s hard to get fully connected bootstrap network, and p-value can’t be found. haven’t test yet, but what if just get rid of the isolate nodes in new method?). This bootstrap method could be extended to other Euclidean models provided that the distribution of distance between nodes given their geodesic distance could be found theoretically or through Monte Carlo method.
However, both permutation and bootstrapping methods are computationally expensive since both classical MDS and hyperbolic MDS need to be applied to each of the permuted or bootstrapping networks.

### v2

@@ 4.2 Network measures
The following network summary measures are used:
@@ 4.3 Simulation details
To evaluate the efficiency of comparing raw stress difference, we tested networks of size 15 to 75 with densities ranging from 0 to 0.9. To compare all three methods, we first tested networks of size 15 to 45 with densities ranging from 0 to 0.9. Then, we focus on large and sparse networks whose sizes ranges from 50 to 200 and densities up to 0.2.
@@ 4.4 Results
4.4.1 Comparing raw stress differences
As shown in Figure 3, we notice that as network size increases, hyperbolic networks tend to be correctly identified. When network size is larger than 60, almost all hyperbolic networks are correctly categorized as hyperbolic. For smaller hyperbolic networks, the probability of correctly identifying hyperbolic geometry gets larger as the density of the network gets larger. However, this method doesn’t work well for Euclidean networks, and larger and/or denser networks tends to be misclassified as hyperbolic network.
Figure 3: Comparing raw stress differences misclassifies almost all Euclidean networks as hyperbolic.
4.4.2 Adding uncertainty from MDS
We compare this approach to the above methods for accounting for uncertainty from MDS: (1) classical permutation tests for MDS and (2) bootstrapped networks from observed geodesic distances. We adopt the language of diagnostic tests. Let sensitivity indicate how well a method correctly identifies hyperbolic networks, and specificity indicate how well a method correctly identifies Euclidean networks.
First, we focus on smaller networks with a wide range of densities in Table 1.
N Density $S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y})$ Permutation Test Bootstrapped Networks
@@ 4.4.2 Adding uncertainty from MDS
Table 1: Smaller Networks
In general, for smaller networks, comparing raw stress differences identifies hyperbolic networks very well but, at the same time, tends to misclassify the majority of Euclidean networks especially as the networks get larger and/or denser. Both the permutation test and bootstrapped approach are more conservative and less likely to falsely categorize Euclidean networks as hyperbolic networks but misclassify the hyperbolic networks more than roughly half the time.
Then to mimic many real world networks, we focused on a particular group of networks - large and sparse networks. In this case, only networks of densities between 0 and 0.2 will be considered.
N $S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y})$ Permutation Test Bootstrapped Networks
@@
Table 2: Large and Sparse Networks
From Table 2, when network size is 100 or larger, comparing raw stress differences categorizes all simulated networks, hyperbolic and Euclidean, as hyperbolic networks and cannot distinguish underlying geometry. Both the permutation test and bootstrapped based approach show an improvement in distinguishing between the two geometries for large and sparse network, which is more common in real-world datasets. The permutation test has better sensitivity, meaning that it is good at correctly identifying hyperbolic networks, while incorporating a bootstrapped network distribution results in better balance between sensitivity and specificity.
@@ 5.1 Karate Club
Papamichalis et al. [2022] use Zachary’s Karate club network as an example. They compare the quality of MDS estimates when the manifold is assumed to be hyperbolic and Euclidean, and conclude that a model with a hyperbolic latent space better suits the data. Using the same package mentioned in this paper, hydra, we find $S_{\mathbb{R}^{2}}(\mathbf{Y})=24.65$ and $S_{\mathbb{H}^{2}}(\mathbf{Y})=18.20$. The stress from hyperbolic space is smaller, which would seem to indicate a higher quality of embedding and that hyperbolic geometry would be better to model the data. Accounting for uncertainty for MDS, we get p-values of 0.1053 from classical MDS permutation tests and 0.1888 from our bootstrap network distribution. This leads us to the conclusion that a Euclidean geometry is better. In Figure 4, we show the distribution of stress differences across the permuted networks (panel a) and bootstrapped networks (panel b).
(a) p-value: 0.1053
(b) p-value: 0.1888
Figure 4: Stress differences across permuted (left panel) and bootstrapped (right panel) networks based on Zachary’s Karate network. All differences are calculated as $S_{\mathbb{H}^{2}}-S_{\mathbb{R}^{2}}$, with a negative difference indicating that hyperbolic space better describes the observed network geodesics. The vertical red lines correspond to the raw stress difference, without accounting for uncertainty.
@@ 5.2 Other Datasets
We have also applied these methods to four additional network datasets [Nepusz et al., 2008, Newman, 2006, Lusseau et al., 2003, Kadushin, 1995]. A comparison of their network characteristics and the results from each of the three different methods is presented below. These networks are relatively sparse and range in size from 28 to 112 nodes.
Across the four networks, three show agreement among all applicable methods on the underlying geometry, and the exception is the French Financial Elite Friendships network (ffe_friend). Similar to Zachary’s Karate network, the raw stress difference method favors a hyperbolic embedding. However, both methods accounting for uncertainty for MDS indicate that a Euclidean geometry provides the better fit. Overall, based on the permutation test and the bootstrapped network methods, UKfaculty and adjnoun are better modeled with hyperbolic geometry, whereas dolphins and ffe_friend are better with Euclidean.
For the UKfaculty network, the bootstrapped network method is not applicable. Several nodes, for example node 73, have very low degree (connected to only two others), and those neighbors are also sparsely connected. This places such nodes far from the rest in latent space, lowers their probability of forming ties, and rarely yields a fully connected graph. Because the bootstrapped network method requires the resampled graph to be fully connected, it is difficult to implement for this dataset.
(a) UKfaculty ($n=81$, dens.=0.178=0.178)
$S_{\mathbb{H}^{2}}-S_{\mathbb{R}^{2}}$ Permutation Bootstrap -19.32 p-value: 0 - $\mathbb{H}^{2}$ $\mathbb{H}^{2}$ -
(b) adjnoun ($n=112$, dens.=0.068=0.068)
$S_{\mathbb{H}^{2}}-S_{\mathbb{R}^{2}}$ Permutation Bootstrap -62.61 p-value: 0 p-value: 0 $\mathbb{H}^{2}$ $\mathbb{H}^{2}$ $\mathbb{H}^{2}$
(c) dolphins ($n=62$, dens.=0.084=0.084)
$S_{\mathbb{H}^{2}}-S_{\mathbb{R}^{2}}$ Permutation Bootstrap 11.51 p-value: 0.99 p-value: 0.99 $\mathbb{R}^{2}$ $\mathbb{R}^{2}$ $\mathbb{R}^{2}$
(d) ffe_friend ($n=28$, dens.=0.175=0.175)
$S_{\mathbb{H}^{2}}-S_{\mathbb{R}^{2}}$ Permutation Bootstrap -3.60 p-value: 0.20 p-value: 0.38 $\mathbb{H}^{2}$ $\mathbb{R}^{2}$ $\mathbb{R}^{2}$
Figure 5: MDS-based classifications of a collection of real world networks. All network plots are created using package igraph, which emphasizes tree-like behavior more common in hyperbolic networks. Each network plot is annotated with its network size, $n$, edge density (dens.), and a table describing the results of the MDS-based classifications.
@@ 6 Discussion
We studied the effectiveness of using observed stress difference from MDS in determining the underlying geometry for latent space models through a simulation study. We have demonstrated that when the stress difference is used alone, the geometry cannot be correctly identified. In fact, comparing the stress difference directly nearly always misclassifies large Euclidean networks as hyperbolic.
Two new methods based on the stress difference from MDS were proposed that account for uncertainty with permutation tests and bootstrapped networks. By studying the same sets of simulated networks, we showed that both methods outperform the original simple stress difference, and work better on large and sparse networks which are commonly seen in real world. However, both permutation and bootstrapping methods are computationally expensive since both classical MDS and hyperbolic MDS need to be applied to each of the permuted or bootstrapping networks.
The advantage of the permutation test is that it retains the density of the observed network and doesn’t rely on the latent space models, but the structure of the network is not considered when generating the permuted networks. On the other hand, the bootstrap method takes the structure of the observed networks into account: if two nodes are currently connected, they are likely closer in the latent space, and thus more likely to be connected in the bootstrapped networks. In this study, we used the Gaussian LPM as the Euclidean latent space model. However, for some networks, $\phi$ and $\tau$ could not be estimated using the ad-hoc method described in Rastelli et al. [2016], thus bootstrapping could not be implemented. Another difficulty in implementing the bootstrap method is that it sometimes produces networks that are not fully connected. As shown above for the UKfaculty network in Section 5.2, bootstrap resampling frequently yielded disconnected networks, consequently, the procedure could not be executed and no p-value could be computed. One potential solution is to remove isolated nodes, but this results in a bootstrapped networks that can have varying network sizes, which is undesirable.
The bootstrap method could be extended to other network models, including more general Euclidean latent space models (i.e., that do not assume a Gaussian link) and latent space models in curved geometries. This requires deriving the model-based conditional distribution of the latent distance between nodes given their geodesic distance. In cases where this can not be found theoretically, Monte Carlo estimation could be used. In the current study, we compare hyperbolic and Euclidean latent space models, using Euclidean geometry as the baseline and hyperbolic geometry in the competing model. A natural extension would be to adopt a spherical geometry as the baseline. Because spherical and hyperbolic spaces are more geometrically distinct than Euclidean and hyperbolic spaces, this choice may further improve the ability to identify networks with underlying hyperbolic structure.